# Fixation and extinction in time-fluctuating spatially structured metapopulations

### v1

@@ Appendix H Asymmetric dichotomous Markov noise & environmental bias
For the sake of simplicity and clarity, in the main text we have focused on symmetric environmental switching. In this section, we relax this assumption and outline how the results of the paper can be generalised to the case when there is an environmental bias, i.e. when there is a different average time spent in the states $\xi=\pm 1$.
Here, we consider the coloured asymmetric dichotomous Markov noise (aDMN), also called telegraph process, $\xi(t)\in\{-1,1\}$ that switches between ±$\pm 1$ according to $\xi\to-\xi$ at rate $\nu_{\pm}$ when $\xi=\pm 1$ [68, 69, 70]. It is convenient to write these asymmetric switching rates as $\nu_{\pm}=\nu(1\mp\delta)$, where $\nu\equiv(\nu_{-}+\nu_{+})/2$ is the mean switching rate , $\delta\equiv(\nu_{-}-\nu_{+})/(\nu_{-}+\nu_{+})=(\nu_{-}-\nu_{+})/(2\nu)$ denotes the switching bias, with $|\delta|\leq 1$ and $\delta>0$ when more time is spent on average in the mild environment [18, 19]. At stationarity, this aDMN has average $\langle\xi(t)\rangle=\delta$ and autocovariance $\langle\xi(t)\xi(t^{\prime})\rangle-\langle\xi(t)\rangle\langle\xi(t^{\prime})\rangle=(1-\delta^{2})e^{-2\nu|t-t^{\prime}|}$ [68, 69, 70, 16, 17].
@@
which is the same expression as Eq. (5), but now driven by the aDMN $\xi(t)$ [16, 17, 18, 19, 21, 22, 32]. The carrying capacity switches back an forth from $K_{+}$ (mild environment, $\xi=1$) to $K_{-}<K_{+}$ (harsh environment, $\xi=-1$) at rate $\nu_{\pm}=\nu(1\mp\delta)$ according to
$K_{+}\xrightleftharpoons[\nu_{-}=\nu(1+\delta)]{\nu_{+}=\nu(1-\delta)}K_{-}.$
@@
Figure 12: Fixation probability $\Phi^{\rm G}$ and mean fixation time $\Theta^{\rm G}$ against switching rate $\nu$ for different values of $\delta$. Red, blue, and yellow represent $\delta=-0.5$, $\delta=0.0$, and $\delta=0.5$, respectively. Markers show simulation results and lines are predictions of Eq. (24). Other parameters are $\Omega=16$, $s=0.01$, $m=10^{-4}$, $K_{+}=200$, and $K_{-}=20$.
Under asymmetric switching, the stationary population distribution of a single deme, both in the absence and presence of migration (see Appendix A.2), is well approximated by stationary density of the piecewise Markov process Eq. (29) ($n$-PDMP), now driven by the aDMN, whose joint density is
$p_{\xi}(\nu,\delta,n)\propto\begin{cases}\frac{1+\delta}{n^{2}}\left(\frac{K_{+}-n}{n}\right)^{\nu(1-\delta)-1}\left(\frac{n-K_{-}}{n}\right)^{\nu(1+\delta)}\quad\text{if $\xi=1$,}\\ \frac{1-\delta}{n^{2}}\left(\frac{K_{+}-n}{n}\right)^{\nu(1-\delta)}\left(\frac{n-K_{-}}{n}\right)^{\nu(1+\delta)-1}\quad\text{if $\xi=-1$},\end{cases}$ (49)
where the proportional factor accounts for the normalization constants. The stationary marginal density of this $n$-PDMP, up to the normalisation constant, is
$p(\nu,\delta,n)$ $=\sum_{\xi}\left(\frac{1+\xi\delta}{2}\right)p_{\xi}(\nu,\delta,n),$ (50)
@@
In the realm of the coarse-grained description discussed in the main text, the regime of weak bottlenecks dominated by the $W/M$ competition can be characterised by the $M$ fixation probability $\Phi^{\rm G}(\nu,\delta,m)$ and unconditional mean fixation time $\Theta^{\rm G}(\nu,\delta,m)$ by Eq. (24) obtained by solving the first-step analysis equations Eq. (21) and Eq. (22) with the transition rates Eq. (19) and Eq. (20) obtained using $N_{\xi}(\nu,\delta)$ averaged over Eq. (50), i.e. $N_{\xi}(\nu,\delta)=\int np_{\xi}(\nu,\delta,n){\rm d}n$. The results of Fig. 12 for a cycle metapopulation show that the predictions coarse-grained description based on the PDMP approximation Eq. (50) are in good agreement with simulation results. $\Phi^{\rm G}(\nu,\delta,m)$ and $\Theta^{\rm G}(\nu,\delta,m)$ are again found to exhibit a non-monotonic dependence on $\nu$, with extrema in the range of intermediate $\nu$. The main effect of $\delta$ is to increase the $M$ fixation probability and reduce the mean fixation time when $\delta>0$, which is intuitively clear since this corresponds to a bias towards the mild state favouring the fixation of $M$.
@@
Figure 13: Near-optimal condition for the idealised treatment on the clique metapopulation. $(\nu,K_{-})$ heatmaps of $\Phi$, $\Theta$, $\Theta_{E}$ and $(1-\Phi)(1-\Theta/\Theta_{E})$ for the $\delta=0.5$ (a-d) and $\delta=-0.5$ (e-h) metapopulations, see Appendix H. Whitespace in panels (c) and (g) indicate the region of the parameter where at least one realisation for those parameters did not reach extinction by $t=10^{5}$. Grey lines in panels (d) and (h) show the near-optimal conditions for the idealised treatment in the asymmetric environment: $\psi(m,K_{-})<1$ below the top horizontal line, $mK_{+}\theta_{E}\frac{1+\delta}{1-\delta}>1$ above the bottom horizontal line, and $\nu(1+\delta)\theta_{E}>1$ above the curved line, while the vertical line indicates where $\nu<1$ and $\theta_{E}$ from Eq. (13). The near-optimal treatment conditions is the yellowish cloud at the centre of the area enclosed by these lines. Other parameters are $\Omega=16$, $m=10^{-4}$, $s=0.1$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@ Appendix I Simulation methods & plots
In this section, we explain how the simulation of the individual-based dynamics of the full model have been implemented. We also outline how we have plotted the simulation/numerical data that we have obtained to produce the figures discussed in the main text and appendices.
In addition to the coarse-grained descriptions of the model that provide us with analytical approximations of metapopulation dynamics in different regimes, we have employed Monte Carlo (MC) methods to simulate the full individual-based model and mirror the dynamics encoded in the ME Eq. (27). In this section, we outline how we have performed the stochastic simulations that we have notably used to test our analytical predictions.
While not exact like other simulation methods (e.g. the Gillespie algorithm [98]), the Monte Carlo algorithm used here improves on computational efficiency making simulations with larger numbers of cells on the metapopulation that run for long times feasible. The questions of computational efficiency and tractability are particular critical in the context of this work in which we study fixation and extinction of spatially arranged populations, a notoriously computationally demanding problem.
@@
In all simulations, each realisation is simulated until fixation, tracking when fixation occurs and which species fixates. In simulations for Figs. 2(e-g), 6, 10, and 13 the simulation then continues until metapopulation extinction or a large fixed time $T$ (here we set $T=10^{5}$). For a given set of parameters, if extinction does not occur in any realisation, the extinction time is not recorded. The data is averaged to obtain the fixation probability, the mean fixation time, and the mean time to extinction where applicable. Furthermore, we record the standard error on the mean for each quantity. In Figs. 2, 4, and 12, $10^{3}$ realisations are ran for each set of parameters, and these data are plotted with the standard error on the mean shows as error bars. In the heatmaps, each data point corresponds to the average value for $10^{3}$ simulations at that point in parameter space. The standard error on the mean is not plotted for the heatmaps.

### v2

@@ Appendix H Asymmetric dichotomous Markov noise & environmental bias
For the sake of simplicity and clarity, in the main text we have focussed on symmetric environmental switching. In this section, we relax this assumption and outline how the results of the paper can be generalised to the case when there is an environmental bias, i.e. when there is a different average time spent in the states $\xi=\pm 1$.
Here, we consider the coloured asymmetric dichotomous Markov noise (aDMN), also called telegraph process, $\xi(t)\in\{-1,1\}$ that switches between $\pm 1$ according to $\xi\to-\xi$ at rate $\nu_{\pm}$ when $\xi=\pm 1$ [68, 69, 70]. It is convenient to write these asymmetric switching rates as $\nu_{\pm}=\nu(1\mp\delta)$, where $\nu\equiv(\nu_{-}+\nu_{+})/2$ is the mean switching rate , $\delta\equiv(\nu_{-}-\nu_{+})/(\nu_{-}+\nu_{+})=(\nu_{-}-\nu_{+})/(2\nu)$ denotes the switching bias, with $|\delta|\leq 1$ and $\delta>0$ when more time is spent on average in the mild environment [18, 19]. At stationarity, this aDMN has average $\langle\xi(t)\rangle=\delta$ and autocovariance $\langle\xi(t)\xi(t^{\prime})\rangle-\langle\xi(t)\rangle\langle\xi(t^{\prime})\rangle=(1-\delta^{2})e^{-2\nu|t-t^{\prime}|}$ [68, 69, 70, 16, 17].
@@
which is the same expression as Eq. (6), but now driven by the aDMN $\xi(t)$ [16, 17, 18, 19, 21, 22, 32]. The carrying capacity switches back and forth from $K_{+}$ (mild environment, $\xi=1$) to $K_{-}<K_{+}$ (harsh environment, $\xi=-1$) at rates $\nu_{\pm}=\nu(1\mp\delta)$ according to
$K_{+}\xrightleftharpoons[\nu_{-}]{\nu_{+}}K_{-}.$
@@
Figure 12: Fixation probability $\Phi^{\rm G}$ and mean fixation time $\Theta^{\rm G}$ against switching rate $\nu$ for different values of $\delta$ for a clique metapopulation. Red, blue, and yellow represent $\delta=-0.5$, $\delta=0.0$, and $\delta=0.5$, respectively. Markers show simulation results and lines are predictions of Eq. (26). Other parameters are $\Omega=16$, $s=0.01$, $m=10^{-4}$, $K_{+}=200$, and $K_{-}=20$.
Under asymmetric switching, the stationary population distribution of a single deme is well approximated by the stationary density of the piecewise Markov process Eq. (30) ($n$-PDMP) [17], now driven by the aDMN, whose joint density is
[…]
where the proportional factor accounts for the normalisation constants. The stationary marginal density of this $n$-PDMP, up to the normalisation constant, is
$p(n;\nu,\delta)$ $=\sum_{\xi}\left(\frac{1+\xi\delta}{2}\right)p_{\xi}(n;\nu,\delta),$ (54)
@@
In the realm of the coarse-grained description discussed in the main text, the regime of weak bottlenecks dominated by the $W/M$ competition can be characterised by the $M$ fixation probability $\Phi^{\rm G}(\nu,\delta,m)$ and unconditional mean fixation time $\Theta^{\rm G}(\nu,\delta,m)$ by Eq. (26) obtained by solving the first-step analysis equations Eq. (23) and Eq. (24) with the transition rates Eq. (21) and Eq. (22) obtained using $N_{\xi}(\nu,\delta)$ averaged over Eq. (54), i.e. $N_{\xi}(\nu,\delta)=\int np_{\xi}(n;\nu,\delta){\rm d}n$. The results of Fig. 12 for a clique metapopulation show that the predictions of the coarse-grained description based on the PDMP approximation Eq. (54) are in good agreement with simulation results. $\Phi^{\rm G}(\nu,\delta,m)$ and $\Theta^{\rm G}(\nu,\delta,m)$ are again found to exhibit a non-monotonic dependence on $\nu$, with extrema in the range of intermediate $\nu$. The main effect of $\delta$ is to increase the $M$ fixation probability and reduce the mean fixation time when $\delta>0$, which is intuitively clear since this corresponds to a bias towards the mild state favouring the fixation of $M$.
@@
Figure 13: Near-optimal condition for the idealised treatment strategy on the clique metapopulation. $(\nu,K_{-})$ heatmaps of $\Phi$, $\Theta$, $\Theta_{E}$ and $(1-\Phi)(1-\Theta/\Theta_{E})$ for the $\delta=0.5$ (a-d) and $\delta=-0.5$ (e-h) metapopulations; see Appendix H. Whitespace in panels (c) and (g) indicate the region of the parameter where at least one realisation for those parameters did not reach extinction by $t=10^{5}$. Grey lines in panels (d) and (h) show the near-optimal conditions for the idealised treatment strategy in the asymmetric environment: $\psi(m,K_{-})<1$ below the top horizontal line, $mK_{+}\theta_{E}\frac{1+\delta}{1-\delta}>1$ above the bottom horizontal line, and $\nu(1+\delta)\theta_{E}>1$ above the curved line, while the vertical line indicates where $\nu<1$ and $\theta_{E}$ from Eq. (15). The near-optimal treatment conditions is the yellowish cloud at the centre of the area enclosed by these lines. Other parameters are $\Omega=16$, $m=10^{-4}$, $s=0.1$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@ Appendix I Simulation methods & plots
In this section, we explain how the simulation of the individual-based dynamics of the full model has been implemented. We also outline how we have plotted the simulation/numerical data that we have obtained to produce the figures discussed in the main text and appendices.
In addition to the coarse-grained descriptions of the model that provide us with analytical approximations of metapopulation dynamics in different regimes, we have employed Monte Carlo (MC) methods to simulate the full individual-based model and mirror the dynamics encoded in the ME of Eq. (28). In this section, we outline how we have performed the stochastic simulations that we have notably used to test our analytical predictions.
While not exact like other simulation methods (e.g. the Gillespie algorithm [97]) due to time-discretisation, the MC algorithm used here improves on computational efficiency making simulations with larger numbers of cells on the metapopulation that run for long times feasible and straightforward to implement. The questions of computational efficiency and tractability are particular critical in the context of this work in which we study fixation and extinction of spatially arranged populations, a notoriously computationally demanding problem. In this context MC algorithms similar to the one used are useful tools to investigate the properties of spatially extended systems; see, e.g., Ref. [98].
@@
In all simulations, each realisation is simulated until fixation, tracking when fixation occurs and which species fixates. In simulations for Figs. 2(e,f), 6, 10, and 13 the simulation then continues until metapopulation extinction or a large fixed time $T$ (here we set $T=10^{5}$). For a given set of parameters, if extinction does not occur in any realisation, the extinction time is not recorded. The data is averaged to obtain the fixation probability, the mean fixation time, and the mean time to extinction where applicable. Furthermore, we record the standard error on the mean for each quantity. In Figs. 2, 4, and 12, $10^{3}$ realisations are ran for each set of parameters, and these data are plotted with the standard error on the mean shows as error bars. In the heatmaps, each data point corresponds to the average value for $10^{3}$ simulations at that point in parameter space. The standard error on the mean is not plotted for the heatmaps.